# PersonalAI: A Systematic Comparison of Knowledge Graph Storage and Retrieval Approaches for Personalized LLM agents

### v2

@@ Abstract
Personalizing language models by effectively incorporating user interaction history remains a central challenge in the development of adaptive AI systems. While large language models (LLMs) combined with Retrieval-Augmented Generation (RAG) have improved factual accuracy, they often lack structured memory and fail to scale in complex, long-term interactions. To address this, we propose a flexible external memory framework based on knowledge graphs, automatically constructed and updated by the LLM itself, and capable of encoding information in multiple formats—including nodes, triplets, higher-order propositions, and episodic traces. Building upon the AriGraph architecture, we introduce a novel hybrid graph design that supports both standard edges and two types of hyperedges, enabling rich and dynamic semantic and temporal representations. Our framework also supports diverse retrieval mechanisms, including A*, water-circle propagation, beam search, and hybrid methods, making it adaptable to different datasets and LLM capacities. We evaluate our system on three benchmarks—TriviaQA, HotpotQA, and DiaASQ—demonstrating that different memory and retrieval configurations yield optimal performance depending on the task. Additionally, we extend the DiaASQ benchmark with temporal annotations and internally contradictory statements, showing that our system remains robust and effective in managing temporal dependencies and context-aware reasoning.
1 Introduction
@@
Recent advances in large language models (LLMs) have sparked growing interest in personalized AI systems capable of adapting to users based on their interaction history. Central to personalization is the challenge of encoding, storing, and retrieving relevant information over long time horizons in a manner that supports efficient reasoning and response generation. While Retrieval-Augmented Generation (RAG) has become a widely used solution, enhancing factual recall by appending retrieved content to prompts, it remains limited by its unstructured nature and weak support for semantic relationships across stored memories.
@@
In this work, we introduce a flexible graph-based memory framework designed to overcome these limitations by enabling structured, customizable representations of long-term memory and supporting advanced reasoning capabilities. Unlike traditional RAG pipelines that rely on dense vector similarity over raw text chunks, our system supports multiple memory formats—nodes, knowledge triplets, higher-order propositions, and episodic traces—and dynamically organizes them into a knowledge graph. This structure allows the agent to represent, update, and access semantic and temporal relationships with far greater control and interpretability.
@@
Equally important, our framework supports a pluggable retrieval interface with multiple reasoning mechanisms, including three variations of A* search, WaterCircle propagation, BeamSearch, and hybrid strategies to adapt retrieval behavior to task demands and model capacity. We demonstrate that different memory and retrieval configurations yield optimal performance on different benchmarks and LLM scales, highlighting the versatility of our approach.
@@
We build our system on top of the AriGraph architecture (Anokhin et al. 2024), originally developed for LLM agents in interactive text environments such as TextWorld. AriGraph continuously maintains a structured knowledge base by extracting triplets from observations, pruning outdated or redundant facts, and integrating episodic and semantic nodes into a unified graph. Our framework extends this foundation by allowing customization of both save (memorization) and search (retrieval) modules, supporting task-specific tuning and component evaluation.
@@
    We present a highly flexible external memory architecture that supports multiple structured formats (nodes, triplets, propositions, episodes) and can be tuned via orthogonal hyperparameters for storage and retrieval.
@@
    We propose and evaluate six retrieval over structured knowledge graphs, achieving superior performance across various datasets compared to Graph-RAG baselines.
@@
2 Related Work
Figure 1: Example of the graph fragment, constructed from natural language text using our method, with simple (green), hyper (yellow) and episodic (blue) nodes
@@ 2 Related Work
In recent years, substantial advancements have been made in open-domain question answering (QA) and the personalization of language models. Techniques leveraging Wikipedia as a broad knowledge source (Chen et al. 2017) have successfully incorporated large-scale machine reading, integrating document retrieval with textual comprehension. Furthermore, dense representation-based methods for passage retrieval (Karpukhin et al. 2020) have demonstrated superior efficacy compared to traditional sparse retrieval approaches, such as TF-IDF and BM25, particularly in scenarios with sufficient training data.
@@
The development of dense retrievers has demonstrated significant progress, particularly through the application of contrastive learning in unsupervised settings, which has shown promising results across diverse scenarios and outperformed conventional approaches such as BM25 (Robertson, Zaragoza et al. 2009) and Contriever (Izacard et al. 2021). Concurrently, pre-trained language models incorporating non-parametric memory access have been proposed for knowledge-intensive tasks. A notable example is retrieval-augmented generation (RAG) models, which integrate parametric and non-parametric memory mechanisms to improve question-answering performance (Lewis et al. 2021).
@@
Recent advancements in unsupervised dense retrieval models, such as ART, have demonstrated the ability to achieve state-of-the-art performance while eliminating reliance on labeled training data (Sachan et al. 2023). In the domain of knowledge graph-based approaches, frameworks such as GraphReader (Li et al. 2024)) incorporate structured reasoning mechanisms to facilitate knowledge extraction and representation, with a particular emphasis on enhancing long-context reasoning capabilities.
@@
For personalized models, AriGraph (Anokhin et al. 2024) introduces a framework that integrates episodic memory and long-term planning using knowledge graphs. Similarly, HippoRAG (Gutiérrez et al. 2024) employs personalized algorithms to improve question-answering (QA) performance by constructing semantic graphs, demonstrating notable advancements over conventional extraction methods. MemWalker (Chen et al. 2023) and PEARL (Sarthi et al. 2024) address challenges associated with context length, proposing architectures capable of efficiently traversing and consolidating information from large-scale documents.
@@
Additionally, ReadAgent (Lee et al. 2024) addresses the challenge of processing long-text contexts by structuring content into memory episodes. Meanwhile, KGP (Wang et al. 2023) proposes Knowledge Graph Prompting, a method that enhances multi-document question answering by constructing knowledge graphs to improve contextual reasoning.
@@
These advancements demonstrate a sustained emphasis on improving language models’ comprehension, retrieval, and personalization capabilities. Such progress facilitates the development of more sophisticated systems, which leverages knowledge graphs to deliver personalized and contextually enriched interactions.
3 Methods
Graph Construction
Figure 2: High level architecture of proposed QA pipeline for generating answers to the questions based on constructed knowledge graph
@@ 3 Methods
In this study, we employ a graph-based knowledge base as external memory to enhance a large language model’s (LLM) question-answering capabilities. To capture and structure information from weakly structured sources, specifically, unstructured natural language texts, we construct a knowledge graph. This graph provides a comprehensive representation of the original text and comprises the following elements:
@@
    Entity nodes, which represent atomic concepts extracted from the text (depicted as green entities in Figure 1);
@@
    Simple relations, which capture direct associations between entities (illustrated by connections between green nodes in Figure 1);
@@
    Propositions, which encapsulate complete atomic thoughts expressed in the text and function as hyperconnections among arbitrary sets of nodes (represented by yellow entities and connections in Figure 1);
@@
    Episodic memories, which preserve the original text passages and serve as hyperconnections linking relevant nodes (denoted by blue entities and connections in Figure 1).
@@
In the terminology defined above, the process of constructing a knowledge graph from weakly structured sources can be decomposed into three key steps: (1) formulating nodes and their relationships, (2) generating propositions using an LLM, and (3) parsing the model’s output to store the extracted information in a structured format (e.g., as subject-relation-object triples). An example prompt for extracting atomic statements and thesis-level claims from textual sources using an LLM is provided in Appendix A.
@@
The identification and retrieval of outdated information within the graph are performed through the following procedure. First, the entities present in the newly extracted triplets and thesis statements are compared against existing nodes in the graph to detect matches. Upon identifying matching nodes, these serve as the initial set for a breadth-first search (BFS), which traverses all associated connections and hyperconnections originating from or terminating at these nodes. Subsequently, a specialized prompt is employed to instruct the LLM to update the retrieved knowledge with the newly extracted data (see Appendix B for the corresponding LLM prompts). If any triplets or thesis statements are successfully updated, the outdated instances are removed from the graph.
Information Searching
Information search pipline on knowledge graph is shown in Figure 2.
@@
As illustrated in Figure 2, the pipeline comprises four primary stages, with the third stage further subdivided into two substages. The pipeline begins by accepting a natural language question as input, which is then processed by the “Query LLM Parser” module to extract key entities. In the second stage, these entities are passed to the “Knowledge Comparator” module, where they are aligned with corresponding nodes in the knowledge graph.
@@
The third stage involves two sequential operations: retrieval and filtering of knowledge graph triplets based on semantic relevance to the input question. First, the “Graph Triplets Retriever” module initiates a graph traversal algorithm, using the matched entities from the second stage as starting vertices to retrieve a set of candidate triplets. Subsequently, the “Triplets Filter” module ranks these triplets by computing their semantic similarity to the question via vector embeddings, retaining only the top $N$ (a predefined hyperparameter) most relevant triplets.
@@
Finally, in the fourth stage, the “QA LLM Generator” module synthesizes a natural language answer conditioned on the filtered triplets. The output of the pipeline is the generated answer, returned as a string.
@@
This pipeline architecture is designed based on three key considerations. First, to achieve an accurate initial approximation of the relevant subgraph, it is essential to align the key entities in the query with semantically similar entities in the existing knowledge base. The reasoning behind this is that the information necessary for generating a correct response is typically localized within the subgraph containing these key entities. Second, the triplets extracted from the knowledge graph exhibit only weak conditioning on the original query, limiting their direct applicability. Third, large language models (LLMs) are constrained by a fixed maximum sequence length for processing inputs in a single inference step, necessitating efficient retrieval and subgraph selection strategies.
Retrieve Algorithms
@@
The primary role of the constructed knowledge graph in our proposed system is to enable the accurate retrieval of information required for responding to specific user queries. This process necessitates a careful optimization between two key criteria: relevance and completeness. Relevance ensures that all retrieved information is directly pertinent to the query, whereas completeness guarantees the inclusion of all necessary contextual data, even if some retrieved elements may be extraneous. Consequently, striking an optimal balance between these factors is essential for achieving efficient and effective knowledge extraction.
@@
To accomplish this, we design and implement advanced retrieval algorithms capable of dynamically balancing the trade-off between completeness and relevance through configurable parameter settings. These algorithms constitute a critical component of the question-answering pipeline, systematically traversing the knowledge graph to aggregate information essential for generating precise and contextually appropriate responses.
@@
A*. The A* algorithm is a widely used method for graph traversal, particularly valued for its ability to efficiently identify shortest paths between nodes in a graph. In the context of our question-answering pipeline, this algorithm extracts and retains triples encountered along these shortest paths while eliminating duplicates based on content. We hypothesize that the triples obtained through this process contain the necessary information for generating accurate responses. For traversal, we treat the graph as unweighted and undirected, using a constant distance metric between adjacent nodes. To optimize pathfinding efficiency, we evaluate three distinct heuristics for the $h$-metric:
@@
    Inner Product (IP): This heuristic computes the dot product of embeddings between the current node and the target node.
@@
    Weighted Shortest Path: This approach scales the inner product metric by the length of the shortest path, determined via Breadth-First Search (BFS).
@@
    Averaged Weighted Shortest Path: This heuristic calculates the average of inner product distances between adjacent nodes along the path from the start to the current node, as well as the direct path from the current node to the target node, and further weights this value by the BFS-derived shortest path length.
WaterCircles. This graph-based extraction method employs a breadth-first search (BFS) algorithm to retrieve relevant knowledge. Query entities are first mapped to their corresponding nodes in the graph. The algorithm begins by exploring adjacent vertices to these entity nodes and iteratively expands to neighboring vertices in subsequent steps, constructing pathways that radiate outward from the initial entities. When pathways originating from different entities intersect, the triplets formed at these intersections are aggregated into a primary list, while all traversed triplets are compiled into a secondary list. The algorithm ultimately returns a subset of triplets, selecting $N$ from the primary list and $K$ from the secondary list, where $N$ and $K$ are configurable hyperparameters.
When the knowledge graph incorporates not only direct entity relations (i.e., “simple” connections) but also associations between entities and text fragments (referred to as “hyper” and “episodic” connections), a modified breadth-first search (BFS) algorithm is employed as follows:
@@
    Traversal Initialization: The search begins at the entities referenced in the query, following the same procedure as in the case of simple entity relations.
@@
    Text Fragment Analysis: During traversal, identified text fragments are examined for occurrences of other query entities distinct from the entity that originated the path to the given fragment. For each fragment, the number of detected entities, denoted as $N_{intersections}$, is computed.
@@
    Ranking Triplets: The list of “hyper” and “episodic” triplets is then sorted in descending order based on $N_{intersections}$.
This triplet extraction strategy aims to enhance relevance and accuracy in retrieving information from the graph-structured knowledge base, thereby bolstering the effectiveness of AI-driven question-answering systems.
BeamSearch. Given a starting node, the algorithm constructs $N$ (a hyperparameter) semantically relevant paths in response to the input query. This approach is inspired by beam search, a token generation strategy commonly employed in large language model (LLM) inference. The resulting paths are consolidated into a single list, with duplicate triplets removed based on their string content. The traversal process is governed by the following hyperparameters:
@@
    Same Path Intersection by Node: If enabled, a path may revisit a node; otherwise, node revisitation is prohibited.
@@
    Diff Paths Intersection by Node: If enabled, distinct paths may share nodes; otherwise, node sharing is disallowed.
@@
    Diff Paths Intersection by Rel: If enabled, distinct paths may share relations (edges); otherwise, relation sharing is forbidden.
@@
Mixed Algorithm. The mixed algorithm integrates the A*, WaterCircles, and BeamSearch strategies to enhance extraction efficacy. By combining these approaches, the algorithm ensures that triplets not captured by one method (e.g., WaterCircles) may still be retrieved by another (e.g., BeamSearch), thereby improving overall recall. The final set of triplets, which is subsequently passed to the LLM for answer generation, constitutes the union of outputs derived from the A*, WaterCircles, and BeamSearch algorithms.
@@
4 Experiment Set-Up
Datasets
The proposed retrieval algorithms are evaluated on the DiaASQ dataset (Li et al. 2023), which comprises user dialogues from a Chinese forum discussing the characteristics of various mobile devices. The dataset also includes structured statements that encapsulate the core semantic content of these dialogues. Leveraging these annotated “true statements”, a set of evaluation questions was procedurally generated to assess the performance of the proposed algorithms.
In addition to the specialized DiaASQ dataset, the HotpotQA and TriviaQA datasets were selected to assess the performance of the proposed QA pipeline. Using these benchmarks, the functionality implemented in the PersonalAI library is systematically compared against existing Retrieval-Augmented Generation (RAG) and GraphRAG approaches.
As a result of preprocessing initial datasets, their subsamples are used to construct knowledge graphs and evaluate proposed pipeline:
  • DiaASQ: number of qa–pairs – 4800; number of relevant contexts – 3483,
  • HotpotQA: number of qa–pairs – 2000; number of relevant contexts – 3933;
  • TriviaQA: number of qa–pairs – 500; number of relevant contexts – 4925.
More details about preprocessing methods and numerical characteristics of resulted datasets can be observed in Appendix E
Models

### v3

@@ Abstract
Personalizing language models that effectively incorporating user interaction history remains a central challenge in development of adaptive AI systems. While large language models (LLMs), combined with Retrieval-Augmented Generation (RAG), have improved factual accuracy, they often lack structured memory and fail to scale in complex, long-term interactions. To address this, we propose a flexible external memory framework based on knowledge graph, which construct and update memory model automatically by LLM itself. Building upon the AriGraph architecture, we introduce a novel hybrid graph design that supports both standard edges and two types of hyper-edges, enabling rich and dynamic semantic and temporal representations. Our framework also supports diverse retrieval mechanisms, including A*, water-circle traversal, beam search and hybrid methods, making it adaptable to different datasets and LLM capacities. We evaluate our system on three benchmarks: TriviaQA, HotpotQA, DiaASQ and demonstrate that different memory and retrieval configurations yield optimal performance depending on the task. Additionally, we extend the DiaASQ benchmark with temporal annotations and internally contradictory statements, showing that our system remains robust and effective in managing temporal dependencies and context-aware reasoning.
Index Terms:
GraphRAG, Graph Traversal Approaches, Knowledge Graphs Generation, MultiAgency, Question Answering,
[…]
=-21pt
I Introduction
@@
TRecent advances in large language models (LLMs) have sparked growing interest in personalized AI systems capable in adapting to users based on their interaction history. Central to personalization is the challenge of encoding, storing and retrieving relevant information over long time horizons in a manner that supports efficient reasoning and response generation. While Retrieval-Augmented Generation (RAG) has become a widely used solution, enhancing factual recall by appending retrieved content to prompts, it remains limited by its unstructured nature and weak support for semantic relationships across stored memories.
@@
In this work, we introduce a flexible graph-based memory framework designed to overcome these limitations by enabling structured, customizable representations of long-term memory and supporting advanced reasoning capabilities. Unlike traditional RAG pipelines that rely on dense vector similarity over raw text chunks, our system supports multiple memory formats: nodes, knowledge triples, thesis statements, episodic traces and dynamically organizes them into a knowledge graph. This structure allows the agent to represent, update and access semantic and temporal relationships with far greater control and interpretability.
@@
Equally important, our framework supports a pluggable retrieval interface with multiple traversal mechanisms, including three variations of A* search, WaterCircle traversal, BeamSearch and hybrid strategies to adapt retrieval behavior to task demands and model capacity. We demonstrate that different memory and retrieval configurations yield optimal performance on different benchmarks and LLM scales, highlighting the versatility of our approach.
@@
We build our system on top of the AriGraph architecture [anokhin2024arigraphlearningknowledgegraph], originally developed for LLM agents in interactive text environments such as TextWorld. AriGraph continuously maintains a structured knowledge base by extracting triples from observations, pruning outdated or redundant facts and integrating episodic and semantic vertices into a unified graph. Our framework extends this foundation by allowing customization of both save (memorization) and search (retrieval) modules, supporting task-specific tuning and component evaluation.
@@
    We present a highly flexible external memory architecture that can be tuned via orthogonal hyperparameters for storage and retrieval.
@@
    We propose and evaluate six retrieval over structured knowledge graph, achieving superior performance across various datasets compared to GraphRAG baselines.
@@
II Related Work
@@
In recent years, substantial advancements have been made in open-domain question answering (QA) and the personalization of language models. Techniques leveraging Wikipedia as a broad knowledge source [chen-etal-2017-reading] have successfully incorporated large-scale machine reading, integrating document retrieval with textual comprehension. Furthermore, dense representation-based methods for passage retrieval [karpukhin-etal-2020-dense] have demonstrated superior efficacy compared to traditional sparse retrieval approaches, such as TF-IDF and BM25, particularly in scenarios with sufficient training data.
@@
The development of dense retrievers has demonstrated significant progress, particularly through integration of contrastive learning with unsupervised settings, which has shown promising results across diverse scenarios and outperformed conventional approaches such as BM25 [robertson2009probabilistic] and Contriever [izacard2021contriever]. Concurrently, pre-trained language models incorporating non-parametric memory access have been proposed for knowledge-intensive tasks. A notable example is retrieval-augmented generation (RAG) models, which integrate parametric and non-parametric memory mechanisms to improve question-answering performance [lewis2021retrievalaugmentedgenerationknowledgeintensivenlp].
@@
Recent advancements in unsupervised dense retrieval models, such as ART, have demonstrated the ability to achieve state-of-the-art performance while eliminating reliance on labeled training data [sachan-etal-2023-questions]. In the domain of knowledge graph-based approaches, frameworks such as GraphReader [li2024graphreaderbuildinggraphbasedagent]) incorporate structured reasoning mechanisms to facilitate knowledge extraction and representation, with a particular emphasis on enhancing long-context reasoning capabilities.
@@
For personalized models, AriGraph [anokhin2024arigraphlearningknowledgegraph] introduces a framework that integrates episodic memory and long-term planning using knowledge graphs. Similarly, HippoRAG [gutiérrez2024hipporagneurobiologicallyinspiredlongterm] employs personalized algorithms to improve question-answering (QA) performance by constructing semantic graphs, demonstrating notable advancements over conventional extraction methods. MemWalker [chen2023walkingmemorymazecontext] and PEARL [sarthi2024raptorrecursiveabstractiveprocessing] address challenges associated with context length, proposing architectures capable to efficiently traversing and consolidating information from large-scale documents.
@@
Additionally, ReadAgent [lee2024humaninspiredreadingagentgist] addresses the challenge of processing long-text contexts by structuring content into memory episodes. Meanwhile, KGP [wang2023knowledgegraphpromptingmultidocument] proposes Knowledge Graph Prompting, a method that enhances multi-document question answering by constructing knowledge graphs to improve contextual reasoning.
@@
These advancements demonstrate a sustained emphasis on improving language models comprehension, retrieval, and personalization capabilities. Such progress facilitates the development of more sophisticated systems, which leverages knowledge graphs to deliver personalized and contextually enriched interactions.
III Methods
III-A Memory Structure
@@
In this study, we employ a graph knowledge base as external memory to enhance a large language model’s (LLM) question-answering capabilities. Memory model $G=(V_{o},E_{o},V_{t},E_{t},V_{e},E_{e})$ consists of semantic $(V_{o},E_{o},V_{t},E_{t})$ and episodic $(V_{e},E_{e})$ memory vertices and edges. In turn, semantic vertices and edges are classified into theses and objects. To capture and structure information from weakly structured sources, specifically, unstructured natural language texts $d_{i}$, the construction of that memory is performing by LLM in automatic manner. This memory graph provides a comprehensive representation of the original texts and comprises the following elements (see Figure 1):
@@
    $V_{o}$ is a set of object vertices. Each object vertex represents atomic concept extracted from corresponding $d_{i}$.
@@
    $E_{o}$ is a set of object edges. Object edge is a tuple $(v,rel,u)$, where $v$, $u$ are object vertices and relrel is a text-attributed relationship between them, that captures its direct associations. Object edges essentially represent triples integrated in the semantic memory.
@@
    $V_{t}$ is a set of thesis vertices. Thesis vertex encapsulate complete atomic thought, expressed in the corresponding $d_{i}$.
@@
    $E_{t}$ is a set of thesis edges. Thesis edges serves as hyper-edges, linking corresponding set of object vertices, extracted from the same $d_{i}$ and belonging to the given $v^{j}_{t}$.
  5. 5.
    $V_{e}$ is a set of episodic vertices. Each episodic vertex corresponds to an original text passage ($v^{i}_{e}=d_{i}$) and serve as hyper-edge, linking corresponding/relative vertices.
  6. 6.
    $E_{e}$ is a set of episodic edges. Each episodic edge $e^{i}_{e}=(v^{i}_{e},V^{i}_{s})$ connects all semantic vertices $V^{i}_{s}$, extracted from $d_{i}$, with each other by the corresponding episodic vertex $v^{i}_{e}$.
Figure 1: Example of a graph fragment, constructed from natural language text using our method, with object (green), thesis (yellow) and episodic (blue) vertices
III-B Memory Construction
@@
By the terminology defined above, the process of constructing a memory graph from weakly structured sources can be decomposed into three key steps: (1) formulating vertices and their edges, (2) generating hyper-edges using an LLM, and (3) parsing the LLM’s output to store extracted information in a structured format (e.g., as subject-relation-object triples). An LLM prompts for extracting semantic memories from textual sources is provided in Appendix A. Full memory construction pipeline (Memorize pipeline) is demonstrated on Figure 2.
Figure 2: High level architecture of proposed Memorize pipeline for LLM-based triples extraction from unstructured texts on natural language and memory construction
@@
The identification and retrieval of outdated information within the memory are performed through the following procedure. First, the vertices present in the newly extracted triples are compared against existing vertices in the memory graph to detect matches. Upon identifying matching vertices, these serve as the initial set for a breadth-first search (BFS), which traverses all associated standart- and hyper-edges. Subsequently, a specialized prompt is employed to instruct LLM to update retrieved knowledge with newly extracted data (see Appendix B for corresponding LLM prompts). If any triple are successfully updated, the corresponding outdated instance are removed from memory model.
III-C Information Search in Memory
Information search pipeline (QA pipeline) for memory model is shown in Figure 3.
Figure 3: High level architecture of proposed QA pipeline for generating answers to the questions based on constructed memory graph
@@ III-C Information Search in Memory
As illustrated in Figure 3, pipeline comprises four primary stages, with the third stage further subdivided into two substages. QA pipeline begins by accepting a natural language question as input, which is then processed by the Entities Extractor module to extract key entities from it. In the second stage, these entities are passed to the Entities to Vertices Matcher module, where they are aligned with corresponding/relevant vertices in the memory graph.
@@
The third stage involves two sequential operations: retrieval and filtering of graph triples based on semantic similarity to the input question. First, the Memory Graph Triples Retriever module initiates a graph traversal algorithm, using the matched entities from the second stage as starting vertices to retrieve a set of candidate triples. Subsequently, the Triples Filter module ranks these triples by computing their semantic similarity to the question via vector embeddings, retaining only the top $N$ (a predefined hyperparameter) most relevant items.
@@
Finally, in the fourth stage, the Conditional Answer Generator module synthesizes a natural language answer, conditioned on a retrieved/filtered triples. The output of QA pipeline is the generated answer in a string format. Used LLM prompts for described modules is provided in Appendix C.
@@
This pipeline architecture is designed based on three key considerations. First, to achieve an accurate initial approximation of the relevant subgraph, it is essential to align the key entities in the input question with semantically similar vertices in the existing memory graph. The reasoning behind this is that the information necessary for generating a correct response is typically localized within the subgraph containing these key entities. Second, the triples extracted from the knowledge graph exhibit only weak conditioning on the input question, limiting their direct applicability. Third, large language models (LLMs) are constrained by a fixed maximum sequence length for processing inputs in a single inference step, necessitating efficient retrieval and subgraph selection strategies.
III-D Retrieval Algorithms
@@
The primary role of the constructed memory graph in our proposed system is to enable the accurate retrieval of information required for responding to specific user questions. This process necessitates a careful optimization between two key criteria: relevance and completeness. Relevance ensures that all retrieved information is directly pertinent to the question, whereas completeness guarantees the inclusion of all necessary contextual data, even if some retrieved elements may be extraneous. Consequently, striking an optimal balance between these factors is essential for achieving efficient and effective knowledge extraction.
@@
To accomplish this, we design and implement advanced retrieval algorithms capable of dynamically balancing the trade-off between completeness and relevance through configurable parameter settings. These algorithms constitute a critical component of the question-answering pipeline, systematically traversing the memory graph to aggregate information essential for generating precise and contextually appropriate responses.
@@
A*. The A* algorithm is a widely used method for graph traversal, particularly valued for its ability to efficiently identify shortest paths between vertices in a graph. In the context of our question-answering pipeline, this algorithm extracts and retains triples encountered along these shortest paths while eliminating duplicates based on a content. We hypothesize that triples, obtained through this process, contains necessary information for generating accurate responses. For traversal, we treat our memory graph as unweighted and undirected, using a constant distance metric between adjacent vertices. To optimize pathfinding efficiency, we evaluate three distinct heuristics for the $h$-metric:
@@
    Inner Product (IP): This heuristic computes dot product of embeddings between current and target vertices.
@@
    Weighted Shortest Path: This approach scales inner product metric by the length of shortest path, determined via Breadth-First Search (BFS).
@@
    Averaged Weighted Shortest Path: This heuristic calculates average of inner product distances between adjacent vertices along the path from start vertex to current vertex, as well as the direct path from current vertex to target vertex, and further weights this value by BFS-derived shortest path length.
WaterCircles. This graph-based extraction method employs a breadth-first search (BFS) algorithm to retrieve relevant knowledge. Query entities are first mapped to their corresponding/relevant vertices in the memory graph. Algorithm begins with exploring adjacent vertices of initial vertices and iteratively expands to neighboring vertices in subsequent steps, constructing radiate pathways. When pathways originating from different starting vertices intersect, triples, formed at these intersections, are aggregated into a primary list, while all traversed triples are compiled into a secondary list. Algorithm ultimately returns a subset of triples, selecting $N$ from the primary list and $K$ from the secondary list, where $N$ and $K$ are configurable hyperparameters.
When the memory graph incorporates not only direct object relations (object triples), but also associations between objects and text fragments (referred to thesis and episodic triples), a modified breadth-first search (BFS) algorithm is employed as follows:
@@
    Traversal Initialization: The search begins at vertices, that matches entities from the input question.
@@
    Text Fragment Analysis: During traversal, identified text fragments are examined for occurrences of other question entities, distinct from the entity that originated the path to the given fragment. For each fragment, the number of detected entities, denoted as $N_{intersections}$, is computed.
@@
    Ranking Triplets: The list of thesis and episodic triplets is then sorted in descending order based on $N_{intersections}$.
This triplet extraction strategy aims to enhance relevance and accuracy in retrieving information from the memory graph, thereby bolstering the effectiveness of AI-driven question-answering systems.
BeamSearch. Given a starting vertex, algorithm constructs $N$ (a hyperparameter) semantically relevant paths in response to the input question. This approach is inspired by beam search, a token generation strategy commonly employed in large language model (LLM) inference. The resulting paths are consolidated into a single list, with duplicate triples removal based on their string content. Traversal process is governed by the following hyperparameters:
@@
    Same Path Intersection by Vertex: If enabled, a path may revisit a vertex; otherwise, vertex revisitation is prohibited.
@@
    Diff Paths Intersection by vertex: If enabled, distinct paths may share vertices; otherwise, vertex sharing is disallowed.
@@
    Diff Paths Intersection by Edge: If enabled, distinct paths may share edges; otherwise, edge sharing is forbidden.
@@
Mixed Algorithm. This algorithm integrates A*, WaterCircles and BeamSearch strategies to enhance extraction efficacy. By combining these approaches, we ensures that triples not captured by one method (e.g., WaterCircles) may still be retrieved by another (e.g., BeamSearch), thereby improving overall recall. The final set of triples, which is subsequently passed to the LLM for answer generation, constitutes the union of outputs derived from A*, WaterCircles and BeamSearch algorithms.
@@
IV Experiment Set-Up
IV-A Datasets
To evaluate proposed retrieval algorithms, we conducted experiments across three distinct benchmarks: DiaASQ [li2023diaasqbenchmarkconversational], HotpotQA and TriviaQA. This selection was designed to evaluate our framework across varying domains, structural complexities and reasoning requirements.
The primary evaluation dataset is DiaASQ, which consists of user dialogues from a Chinese forum focused on mobile device characteristics. A key feature of this dataset is the inclusion of structured ”true statements” that encapsulate the core semantic content of each dialogue. We procedurally generated evaluation questions from these statements to ensure precise assessment. To further evaluate framework’s capability in handling temporal dynamics and contradictory information, critical for personalized agents, we extended DiaASQ with explicit temporal annotations and internally contradictory statements.
To ensure broad applicability and mitigate potential dataset bias and limited domain diversity, we supplemented DiaASQ with two widely-used, general-domain QA benchmarks:
  1. 1.
    HotpotQA: Selected for its requirement of multi-hop reasoning across multiple documents.
  2. 2.
    TriviaQA: Chosen for its factoid-style, open-domain questions that test broad knowledge retrieval.
These datasets provide complementary challenges, moving beyond a single domain (mobile devices) to perform evaluation on general world knowledge and complex reasoning.
Considering the computational and engineering complexity of constructing and traversing large memory graphs, we created manageable yet representative subsets from the original datasets. This step was necessary to enable the iterative experimentation required for tuning multiple retrieval algorithms and LLM configurations within practical resource constraints. The preprocessing involved filtering contexts by length and, for TriviaQA, segmenting long documents into coherent chunks. The resulting subsets used for knowledge graph construction and QA pipeline evaluation are summarized in Table I. Detailed preprocessing steps and dataset statistics are provided in Appendix D.
Dataset #qa-pairs #contexts
DiaASQ 4800 3483
HotpotQA 2000 3933
TriviaQA 500 4925
TABLE I: Characteristics of prepared datasets for QA pipeline evaluation
Using these benchmarks, memorization and question-answering functionality of our PersonalAI framework is systematically compared against existing Retrieval-Augmented Generation (RAG) and GraphRAG baselines, as detailed in subsequent sections.
IV-B Models